How do you categorize the RSA Algorithm? Assuming p and q are given, write the pseudo code for encryption and decryption.
Answer #1

Step 1:-select p,q two prime no.

Step2:-n=p*q

Φ(n)=(p-1)*(q-1)

step 3:-find public key e

gcd(e,phi(n))=1 and 1<e<phi(n)

step 4:- find private key d

d *e mod phi(n)=1

step 5:-encrypt the messy m

c=m^e mod n

step 6 decrypt

m=c^d mod n